# Bayesian Self-Distillation for Image Classification

### v1

@@ 3.2 Bayesian Self-Distillation
Model: Neural network $f$ with parameters $\bm{\theta}$, optimizer $h$
@@
   $\hat{\mathbf{y}}_{i}^{t}\leftarrow f(\mathbf{x}_{i};\bm{\theta}),\quad\forall i\in B$ $\triangleright$ Model prediction
   $\mathcal{L}\leftarrow\frac{1}{|B|k}\sum_{i\in B}\ell\big(\hat{\mathbf{y}}_{i}^{t},\mathbf{y}_{i}^{t-1}\big)$ $\triangleright$ Compute loss
   Compute $\nabla_{\bm{\theta}}\mathcal{L}$
   $\theta\leftarrow h\left(\bm{\theta},\nabla_{\bm{\theta}}\mathcal{L}\right)$ $\triangleright$ Gradient descent step
   $\mathbf{y}^{t}_{i}\leftarrow\frac{\gamma A_{i}^{t-1}}{\gamma A_{i}^{t-1}+1}\mathbf{y}_{i}^{t-1}+\left(1-\frac{\gamma A_{i}^{t-1}}{\gamma A_{i}^{t-1}+1}\right)\hat{\mathbf{y}}_{i}^{t},\quad\forall i\in B$ $\triangleright$ Update targets
   $A_{i}^{t}=\gamma A_{i}^{t-1}+1,\quad\forall i\in B$ $\triangleright$ Update prior hyperparameters
@@
Return $\bm{\theta}$ $\triangleright$ Trained parameters
3.3 Relationship between BSD and other methods
Similarly,
@@ 3.3 Relationship between BSD and other methods
i.e. the weight of new observations converges to $1-\gamma$ exponentially. In the case of non-zero $\epsilon$, and if we initialize at the fixed point $A_{i}^{0}=\frac{1}{1-\gamma}$ for $\gamma$ close to $1$, BSD will approximate label smoothing (szegedy2016rethinking).
Related self-distillation methods can be interpreted as special cases of BSD. Conventional training corresponds to taking $c\to\infty$, which fixes the target distribution and prevents Bayesian updating. In this limit, the framework reduces to standard distribution matching, as KL-divergence and cross-entropy are equivalent under fixed one-hot targets. PS-KD (kim2021self) is recovered by setting $\gamma=0$, so that only the most recent prediction contributes to the target. DLB (shen2022self) performs the same Bayesian update as PS-KD but applies it at the mini-batch level rather than epoch-wise. Meanwhile, TE (laine2016temporal) corresponds to assuming an improper prior (a zero vector) and applying an explicit weighting schedule to the accumulated evidence. Although these methods differ from BSD in that they include a supervised loss term based on one-hot labels, apply temperature scaling or modify the loss function, all of them implicitly construct a target distribution from an accumulation of past predictions. BSD makes this shared structure explicit by interpreting it as Bayesian evidence aggregation under a Dirichlet prior.
@@ 4 Experiments
Experimental Setup. We evaluate ResNet (he2016deep), DenseNet (huang2017densely), and ViT (dosovitskiy2020image) models on CIFAR-10 (krizhevsky2009learning), CIFAR-100 (krizhevsky2009learning), Tiny ImageNet (tinyimagenet) and ImageNet russakovsky2015imagenet. We report the average of three runs. For BSD, we set $\epsilon=0$, $\gamma=0.95$, and use $c=1000$ and $c=50$ for the CNNs and ViTs, respectively, on all datasets except for ImageNet. On Imagenet, we set $\epsilon=0.05$, $\gamma=0.99$, and use $c=400$. The experimental setup is described in more detail in the appendix, including the hyperparameters used for the baselines.
4.1 Emergence of Dark Knowledge
We first validate that dark knowledge emerges naturally during training, and show that BSD promotes further discovery of underlying structures in the data. In an attempt to formalize the notion of dark knowledge, we decompose the output of the network, for an input $\mathbf{x}_{i}$ with integer label $\mathrm{y}_{i}=\underset{k}{\mathrm{argmax}}\ \mathbf{y}_{i,k}^{0}$, as
$f(\mathbf{x}_{i})=\bm{\mu}_{y_{i}}+\delta(\mathbf{x}_{i}).$ (9)
(a) Conventional training
(b) BSD
Figure 2: Inter-class component $\bm{\mu}$ of dark knowledge. Semantical patterns emerge between classes, accentuated by BSD (ResNet-18, CIFAR-10).
(a) Conventional training
(b) BSD
Figure 3: Sample-specific component of dark knowledge. BSD promotes learning of sample-specific information (ResNet-18, CIFAR-10).
Here, $\bm{\mu}_{y_{i}}\in\Delta^{K}$ is a row of the matrix with elements
$\mu_{ij}=\frac{1}{2}\left(\mathbb{E}[f(X)\mid Y=i]_{j}+\mathbb{E}[f(X)\mid Y=j]_{i}\right)$ (10)
and captures the inter-class component of dark knowledge under a symmetry constraint. The function $\delta:\mathbb{R}^{d}\to\mathbb{R}^{K}$ captures the sample-specific deviation from $\bm{\mu}_{y_{i}}$ and is calculated as $\delta(\mathbf{x}_{i})=f(\mathbf{x}_{i})-\bm{\mu}_{y}$.
We visualize $\bm{\mu}$ and $\delta$ from a ResNet-18 trained on CIFAR-10 in Figures 2 and 3. The log-scale heatmaps of $\bm{\mu}$ in Figure 2 reveal that the networks learn meaningful inter-class relationships. For instance, they capture similarities between animals but also between bird and airplane, with these clusters being more prominent and the probabilities larger for BSD. The per-sample absolute deviation from $\bm{\mu}$ presented in Figure 3 validates the emergence of $\delta$ and shows that BSD promotes learning of sample-specific information.
@@ 4.2 Classification Results
Methods compared. We benchmark BSD against conventional training (baseline), Temporal Ensembling (TE) (laine2016temporal), self-Distillation from Last mini-Batch (DLB) (shen2022self) and Progressive Self-Knowledge Distillation (PS-KD) (kim2021self). We do not compare with methods such as (zhang2019your; zhang2022self), as our focus is on modifying label distributions rather than model architectures. An ensemble of three models is included for reference, whose knowledge is also distilled into a single model of the same architecture, for a comparison with conventional knowledge distillation.
@@ 4.2.1 Generalization
The main results are reported in Table 1. In all experiments, BSD improves test accuracy relative to both the baseline and related methods, surpassing conventional knowledge distillation and approaching ensemble performance.
The performance gains are most pronounced on CIFAR-100 and TinyImageNet. For ResNet and DenseNet, BSD boosts accuracy by about 3 percentage points (pp) over the baseline and by more than 1 pp over the strongest related methods on both datasets. In contrast, on CIFAR-10, where baseline performance is already high, the improvements are more modest. Similarly, for ViT-B/16, we observe small but consistent gains across all datasets. The trend continues on ImageNet, where BSD improves accuracy by roughly 1 pp over the baseline and 0.5 pp over the best-performing related method, indicating that it also scales effectively to more complex datasets.
Table 1: Test set accuracy on CIFAR-10, CIFAR-100, and TinyImageNet. The models are trained using standard training (baseline), related methods (laine2016temporal; shen2022self; kim2021self), and the proposed method (BSD). The best results are highlighted in bold.
Dataset Model Baseline (%) TE (%) DLB (%) PS-KD (%) BSD (%) KD (%) Ens. (%)
@@
ImageNet ResNet-152 $78.55{\scriptstyle\pm 0.05}$ $8.45{\scriptstyle\pm 0.07} 77$$7.46{\scriptstyle\pm 0.14}$ $78.99{\scriptstyle\pm 0.10}$ $\mathbf{79.47}\,{\scriptscriptstyle{\pm 0.07}}$ $79.37{\scriptstyle\pm 0.04}$ $80.24$
Table 2: ECE and NLL on CIFAR-100. The models are trained using standard training (baseline), related methods (laine2016temporal; shen2022self; kim2021self), and the proposed method (BSD). The best results are highlighted in bold.
Architecture Metric Baseline TE DLB PS-KD BSD
@@
4.2.2 Calibration
Given the importance of calibrated probability estimates in many tasks, we evaluate the calibration of models trained on CIFAR-100 using Expected Calibration Error (ECE) (naeini2015obtaining) and Negative Log Likelihood (NLL). The results are included in Table 2, where BSD demonstrates superior calibration when compared to related methods (DLB). Figure 4 includes reliability diagrams where BSD’s curves lie closest to the diagonal line, indicating better calibration. Furthermore, we compare BSD with standard and distillation-based calibration methods in Table 7 (appendix), where it achieves state-of-the-art results across all metrics.
The improvements are significant compared to the baseline, with BSD reducing ECE by more than 60% and NLL by over 70% for the convolutional networks. Relative to the best-performing related method, BSD lowers ECE by nearly 40% and NLL by about 29%. For ViT-B/16, the gains are more modest, with a 22% reduction in ECE and a 25% reduction in NLL over the baseline, alongside small but consistent improvements over related methods. As discussed in Section 4.2.5, careful hyperparameter selection can further improve BSD’s calibration, e.g. an ECE of 1.33% for ResNet-50 on CIFAR-100.
@@ 4.2.3 Robustness
To assess our method’s performance under less ideal conditions, we evaluate the models trained on CIFAR-10 on corrupted and perturbed images, and introduce symmetric and asymmetric label noise.
@@
Figure 4: Reliability Diagrams for CIFAR-100 Models. A curve closer to the diagonal indicates better calibration. The models are trained using standard training (baseline), related methods (laine2016temporal; shen2022self; kim2021self), and the proposed method (BSD).
Corruptions and perturbations. We evaluate the robustness to corruptions and perturbations of BSD and related methods by evaluating the models trained on CIFAR-10 on CIFAR-10-C and CIFAR-10-P (hendrycks2019benchmarking). The results are included in Table 3, where we report test set accuracy and mean Flip Probability (mFP) for CIFAR-10-C and CIFAR-10-P, respectively. BSD yields the highest accuracy and the lowest mFP for all models.
On CIFAR-10-C, BSD improves robustness across all architectures, raising accuracy by about $1.5$ pp for both ResNet-18 and DenseNet-121 and by $0.4$ pp for ViT-B over the baseline, slightly surpassing the best related methods. On CIFAR-10-P, BSD achieves the lowest mean flip probability for all models, reducing mFP by around $10\%$ compared to the best-performing related methods.
Table 3: Performance on CIFAR-10-C and CIFAR-10-P. Accuracy and Mean Flip Probability of BSD, baseline, and related methods (laine2016temporal; shen2022self; kim2021self) under corruptions and perturbations. The best results for each metric are highlighted in bold.
@@
Label noise. We generate symmetric label noise by randomly reassigning a percentage of all labels, and asymmetric noise using the class-dependent definition from prior work (tanaka2018joint). Figure 5 shows test error over epochs for the previously mentioned self-distillation methods under 20% label noise for ResNet-18 trained on CIFAR-10. BSD and TE are obtain the highest accuracy under label noise, while the methods that rely more on the hard targets perform worse. Notably, BSD flattens the characteristic double descent curve observed for other methods, indicating a more robust learning process.
We compare BSD with lightweight regularization methods under noisy labels, including Label Smoothing (LS) (szegedy2016rethinking), Symmetric Cross Entropy Learning (SL) (wang2019symmetric), MixUp (zhang2017mixup) and Temporal Ensembling (TE) laine2016temporal. We report the average best obtained accuracy, which for BSD is typically close to the final accuracy, but diverges in varying amounts for the remaining methods. The results are included in upper section of Table 4, where BSD yields the highest accuracy for all noise levels except under $10\%$ asymmetric noise. We hypothesize that the strong performance of BSD is due to the complete detachment from the original targets during training, which would explain the robustness under high noise levels.
To compare BSD with state-of-the-art methods for learning with noisy labels, we combine it with techniques from self-supervised learning (for more information refer to the Appendix). We restrict our analysis to single-stage, single-network approaches, as multi-stage or ensemble methods may be used to further enhance BSD’s performance. We include OT-Filter (feng2023ot), Curriculum and Structure-aware Optimal Transport (CSOT) (chang2023csot) and Dynamic and Uniform Label Correction (DULC) (xu2025revisiting). Like other works, we consider a PreAct ResNet-18 (he2016identity), trained with SGD for 300 epochs with a batch size of $128$ and weight decay of $5e^{-4}$, learning rate of 0.02 and report the average best obtained accuracy. Results are summarized in the lower section of Table 4, where BSD consistently matches or outperforms existing state-of-the-art methods.
4.2.4 Augmentation
As data augmentation is a standard component in image classification, we examine the interaction of BSD with two widely used augmentation methods, CutOut (devries2017improved) and CutMix (yun2019cutmix). CutOut masks out a patch of an image with zeros, while CutMix replaces the patch with a segment from another image and interpolates the labels. For BSD, the augmentations are applied to half of the images in each mini-batch, while the remaining half is used for distillation. This reduces the frequency of label updates but also changes CutMix to combine soft targets rather than one-hot labels. To compensate, we set $\gamma=0.9$ and $c=500$. We include DLB (shen2022self) and conventional training for comparison. Results are reported in Table 5.
Table 4: Test set accuracy under symmetric and asymmetric label noise. Performance of lightweight methods (top, ResNet-18) (szegedy2016rethinking; wang2019symmetric; zhang2017mixup; laine2016temporal) and state-of-the-art methods (bottom, PreAct ResNet-18) (feng2023ot; chang2023csot; xu2025revisiting) on CIFAR-10. The best results are highlighted in bold.
@@ 4.2.4 Augmentation
BSD $93.39{\scriptstyle\pm 0.05}$ $90.71{\scriptstyle\pm 0.26}$ $77.50{\scriptstyle\pm 0.49}$ $94.18{\scriptstyle\pm 0.07}$ $93.20{\scriptstyle\pm 0.14}$ $91.65{\scriptstyle\pm 0.06}$
@@
BSD+ $\mathbf{96.9}\,{\scriptscriptstyle\pm 0.04}$ $\mathbf{96.2}\,{\scriptscriptstyle\pm 0.06}$ $\mathbf{95.1}\,{\scriptscriptstyle\pm 0.06}$ $\mathbf{97.0}\,{\scriptscriptstyle\pm 0.12}$ $\mathbf{96.3}\,{\scriptscriptstyle\pm 0.13}$ $\mathbf{95.6}\,{\scriptscriptstyle\pm 0.25}$
Figure 5: Test Error Over Epochs Under 20% Label Noise on for ResNet-18 on CIFAR-10. The models are trained using standard training (baseline), related methods (laine2016temporal; shen2022self; kim2021self), and the proposed method (BSD).
Table 5: Accuracy on CIFAR-10, CIFAR-100, and TinyImageNet with CutMix and CutOut. The models are trained using standard training (baseline), DLB (shen2022self) and the proposed method (BSD). The best results are highlighted in bold.

### v2

@@ 3.3 Relationship between DPS and other methods
Model: Neural network $f$ with parameters $\bm{\theta}^{0}$, optimizer $h$, loss function $\ell$
@@
   $\hat{\mathbf{y}}_{i}^{t}\leftarrow f(\mathbf{x}_{i};\bm{\theta}),\quad\forall i\in B$
   $\mathcal{L}\leftarrow\frac{1}{|B|}\sum_{i\in B}\ell\big(\hat{\mathbf{y}}_{i}^{t},\mathbf{y}_{i}^{t-1}\big)$
   $\bm{\theta}^{t}\leftarrow h\left(\bm{\theta}^{t-1},\nabla_{\bm{\theta}^{t-1}}\mathcal{L}\right)$
   $\mathbf{y}^{t}_{i}\leftarrow\frac{\gamma A_{i}^{t-1}}{\gamma A_{i}^{t-1}+1}\mathbf{y}_{i}^{t-1}+\left(1-\frac{\gamma A_{i}^{t-1}}{\gamma A_{i}^{t-1}+1}\right)\hat{\mathbf{y}}_{i}^{t}$
   $A_{i}^{t}\leftarrow\gamma A_{i}^{t-1}+1,\quad\forall i\in B$
@@
Return $\bm{\theta}^{t}$
Furthermore,
@@
i.e. the weight of new observations converges to $1-\gamma$ exponentially. In the case of non-zero $\epsilon$, and if we initialize at the fixed point $A_{i}^{0}=\frac{1}{1-\gamma}$ for $\gamma$ close to 11, DPS approximates label smoothing (Szegedy et al., 2016).
Similarly, PS-KD (Kim et al., 2021) is recovered by setting $\gamma=0$, so that only the most recent prediction contributes to the target. DLB (Shen et al., 2022) performs the same Bayesian update as PS-KD but applies it at the mini-batch level rather than epoch-wise. Meanwhile, TE (Laine and Aila, 2016) corresponds to assuming an improper prior (a zero vector) and applying an explicit weighting schedule to the accumulated evidence. Although these methods differ from DPS in that they include a supervised loss term based on one-hot labels, apply temperature scaling or modify the loss function, all of them implicitly construct a target distribution from an accumulation of past predictions. DPS makes this shared structure explicit by interpreting it as Bayesian evidence aggregation under a Dirichlet prior.
@@ 4 Experiments
Experimental Setup. We evaluate ResNet (He et al., 2016a), DenseNet (Huang et al., 2017), and ViT (Dosovitskiy et al., 2020) models on CIFAR-10 (Krizhevsky et al., 2009), CIFAR-100 (Krizhevsky et al., 2009), Tiny ImageNet (Stanford CS231n, 2017) and ImageNet Russakovsky et al. (2015). We report the average of three runs. For DPS, we set $\epsilon=0$, $\gamma=0.95$, and use $c=1000$ and $c=50$ for the CNNs and ViTs, respectively, on all datasets except for ImageNet. On ImageNet, we set $\epsilon=0.05$, $\gamma=0.99$, and use $c=1000$. The experimental setup is described in more detail in the appendix, including the hyperparameters used for the baselines.
@@ 4.1 Classification Results
Methods compared. We benchmark DPS against conventional training (baseline), Temporal Ensembling (TE) (Laine and Aila, 2016), self-Distillation from Last mini-Batch (DLB) (Shen et al., 2022) and Progressive Self-Knowledge Distillation (PS-KD) (Kim et al., 2021). We do not compare with methods such as (Zhang et al., 2019, 2021), as our focus is on modifying label distributions rather than model architectures. An ensemble of three models is included for reference, whose knowledge is also distilled into a single model of the same architecture, for a comparison with conventional knowledge distillation.
@@ 4.1.1 Generalization
The main results are reported in Table 1. In all experiments, DPS improves test accuracy relative to both the baseline and related methods, surpassing conventional knowledge distillation and approaching ensemble performance.
The performance gains are most pronounced on CIFAR-100 and TinyImageNet. For ResNet and DenseNet, DPS boosts accuracy by about 3 percentage points (pp) over the baseline and by more than 1 pp over the best-performing related methods on both datasets. On CIFAR-10, where baseline performance is high, the improvements are more modest. Similarly, for ViT-B/16, we observe small but consistent gains across all datasets. On ImageNet, DPS improves accuracy for ResNet-152 by about 1.5 and 0.7 pp, and for DenseNet-264 by about 2.2 and 1.0 pp, over the baseline and best-performing related method, respectively.
Table 1: Test set accuracy on CIFAR-10, CIFAR-100, TinyImageNet and ImageNet. The models are trained using standard training (baseline), related methods (Laine and Aila, 2016; Shen et al., 2022; Kim et al., 2021), and the proposed method (DPS). The best results are highlighted in bold.
Dataset Model Baseline (%) TE (%) DLB (%) PS-KD (%) DPS (%) KD (%) Ens. (%)
@@
ImageNet ResNet-152 $78.40{\scriptstyle\pm 0.16}$ $78.46{\scriptstyle\pm 0.10}$ $78.54{\scriptstyle\pm 0.14}$ $79.18{\scriptstyle\pm 0.06}$ $\mathbf{79.88}\,{\scriptscriptstyle{\pm 0.11}}$ $79.73{\scriptstyle\pm 0.05}$ $80.35$
DenseNet-264 $76.55{\scriptstyle\pm 0.02}$ $6.57{\scriptstyle\pm 0.07} 77$$20{\scriptstyle\pm 0.07} 77.$$7.83{\scriptstyle\pm 0.02}$ $\mathbf{78.79}\,{\scriptscriptstyle{\pm 0.12}}$ $78.56{\scriptstyle\pm 0.06}$ $79.69$
4.1.2 Calibration
Given the importance of calibrated probability estimates in many tasks, we evaluate the calibration of models trained on CIFAR-100 using Expected Calibration Error (ECE) (Naeini et al., 2015) and Negative Log Likelihood (NLL). The results are included in Table 4.1.2, where DPS demonstrates superior calibration compared to related methods. Figure 2 includes reliability diagrams where DPS’s curves lie closest to the diagonal line, indicating better calibration. Furthermore, we compare our framework with more calibration methods in Table 6 (Appendix), where DPS achieves state-of-the-art calibration across all metrics.
Table 2: ECE and NLL on CIFAR-100. The models are trained using standard training (baseline), related methods (Laine and Aila, 2016; Shen et al., 2022; Kim et al., 2021), and the proposed method (DPS). The best results are highlighted in bold.
Architecture Metric Baseline TE DLB PS-KD DPS
@@ 4.1.2 Calibration
The improvements are significant compared to the baseline, with DPS reducing ECE by more than 60% and NLL by over 70% for the convolutional networks. Relative to the best-performing related method, DPS lowers ECE by nearly 40% and NLL by about 29%. For ViT-B/16, the gains are more modest, with a 22% reduction in ECE and a 25% reduction in NLL over the baseline, alongside small but consistent improvements over related methods. As discussed in Section 4.3, careful hyperparameter selection can further improve DPS’s calibration, e.g. reaching an ECE of 1.33% for ResNet-50 on CIFAR-100.
@@ 4.1.3 Robustness
To assess our framework’s performance under more realistic conditions, we evaluate the models trained on CIFAR-10 on corrupted and perturbed images, and introduce symmetric and asymmetric label noise.
Corruptions and perturbations. We evaluate the robustness to corruptions and perturbations of DPS and related methods by evaluating the models trained on CIFAR-10 on CIFAR-10-C and CIFAR-10-P (Hendrycks and Dietterich, 2019). The results are included in Table 3, where we report test set accuracy and mean Flip Probability (mFP) for CIFAR-10-C and CIFAR-10-P, respectively. DPS yields the highest accuracy and the lowest mFP for all models.
On CIFAR-10-C, DPS improves robustness across all architectures, raising accuracy by about $1.5$ pp for both ResNet-18 and DenseNet-121 and by $0.4$ pp for ViT-B over the baseline, slightly surpassing the best related methods. On CIFAR-10-P, DPS achieves the lowest mean flip probability for all models, reducing mFP by around $10\%$ compared to the best-performing related methods.
@@
Figure 2: Reliability Diagrams for CIFAR-100 Models. A curve closer to the diagonal indicates better calibration. The models are trained using standard training (baseline), related methods (Laine and Aila, 2016; Shen et al., 2022; Kim et al., 2021), and the proposed method (DPS).
Label noise. We generate symmetric label noise by randomly reassigning a percentage of all labels, and asymmetric noise using the class-dependent definition from prior work (Tanaka et al., 2018). Figure 4 shows test error over epochs for the previously mentioned self-distillation methods under 20% label noise for ResNet-18 trained on CIFAR-10. DPS and TE obtain the highest accuracy under label noise, while the methods that rely more on the hard targets perform worse. Notably, DPS flattens the characteristic double descent curve observed for other methods, indicating a more robust learning process.
Table 3: Performance on CIFAR-10-C and CIFAR-10-P. Accuracy and Mean Flip Probability of DPS, baseline, and related methods (Laine and Aila, 2016; Shen et al., 2022; Kim et al., 2021) under corruptions and perturbations. The best results for each metric are highlighted in bold.
@@
Figure 3: Test set accuracy under symmetric and asymmetric label noise. Performance of lightweight methods (top, ResNet-18) (Szegedy et al., 2016; Wang et al., 2019; Zhang et al., 2017; Laine and Aila, 2016) and state-of-the-art methods (bottom, PreAct ResNet-18) (Feng et al., 2023; Chang et al., 2023; Xu et al., 2025) on CIFAR-10. The best results are highlighted in bold.
@@
DPS $93.39{\scriptstyle\pm 0.05}$ $90.71{\scriptstyle\pm 0.26}$ $77.50{\scriptstyle\pm 0.49}$ $94.18{\scriptstyle\pm 0.07}$ $93.20{\scriptstyle\pm 0.14}$ $91.65{\scriptstyle\pm 0.06}$
@@
DPS+ $\mathbf{96.9}\,{\scriptscriptstyle\pm 0.04}$ $\mathbf{96.2}\,{\scriptscriptstyle\pm 0.06}$ $\mathbf{95.1}\,{\scriptscriptstyle\pm 0.06}$ $\mathbf{97.0}\,{\scriptscriptstyle\pm 0.12}$ $\mathbf{96.3}\,{\scriptscriptstyle\pm 0.13}$ $\mathbf{95.6}\,{\scriptscriptstyle\pm 0.25}$
We compare DPS with lightweight regularization methods under noisy labels, including Label Smoothing (LS) (Szegedy et al., 2016), Symmetric Cross Entropy Learning (SL) (Wang et al., 2019), MixUp (Zhang et al., 2017) and Temporal Ensembling (TE) Laine and Aila (2016). We report the average best obtained accuracy, which for DPS is typically close to the final accuracy, but diverges in varying amounts for the remaining methods. The results are included in upper section of Table 3, where DPS yields the highest accuracy for all noise levels except under $10\%$ asymmetric noise. We hypothesize that the strong performance of DPS is due to the complete detachment from the original targets during training, which would explain the robustness under high noise levels.
To compare DPS with state-of-the-art methods for learning with noisy labels, we combine it with techniques from self-supervised learning (for more information refer to the Appendix). We restrict our analysis to single-stage, single-network approaches, as multi-stage or ensemble methods may be used to further enhance DPS’s performance. We include OT-Filter (Feng et al., 2023), Curriculum and Structure-aware Optimal Transport (CSOT) (Chang et al., 2023) and Dynamic and Uniform Label Correction (DULC) (Xu et al., 2025). Like other works, we consider a PreAct ResNet-18 (He et al., 2016b), trained with SGD for 300 epochs with a batch size of $128$, weight decay of $5e^{-4}$, and learning rate of 0.02, while reporting the average best obtained accuracy. Results are summarized in the lower section of Table 3, where DPS consistently matches or outperforms existing state-of-the-art methods.
Figure 4: Test Error Over Epochs Under 20% Label Noise for ResNet-18 on CIFAR-10. The models are trained using standard training (baseline), related methods (Laine and Aila, 2016; Shen et al., 2022; Kim et al., 2021), and the proposed method (DPS).
4.1.4 Augmentation
As data augmentation is a standard component in image classification pipelines, we examine the interaction of DPS with two widely used augmentation methods: CutOut (DeVries and Taylor, 2017) and CutMix (Yun et al., 2019). CutOut masks out a patch of an image with zeros, while CutMix replaces a patch with a segment from another image and interpolates the labels. For DPS, the augmentations are applied to half of the images in each mini-batch, while the remainder is used for distillation. To compensate for the reduced frequency of label updates, we reduce $c$ by half (except for ImageNet, where we keep it at $c=1000$), while doubling the weight of new predictions ($1-\gamma$). Results are reported in Table 4, where we include DLB (Shen et al., 2022) and conventional training for comparison.
Table 4: Accuracy on CIFAR-10, CIFAR-100, and TinyImageNet with CutMix and CutOut. The models are trained using standard training (baseline), DLB (Shen et al., 2022) and the proposed method (DPS). The best results are highlighted in bold.